# Modular invariance groups and defect McKay-Thompson series

### v1

@@ 2 TDLs from monstralizers
Our next pair of monstralizer VOAs are the $\mathbb{Z}_{4}$ parafermion CFT and its commutant in $V^{\natural}$. In this case, the Verlinde lines do exhaust the fusion category, and are given along with their eigenvalues in table 3 [11].
$(l,m)$ $(4,4)$ $(4,2)$ $(4,0)$ $(4,-2)$ $(3,3)$ $(3,1)$ $(3,-1)$ $(1,1)$ $(2,0)$ $(2,2)$
$\theta$ $1$ $-1$ $1$ $-1$ $i$ $-i$ $i$ $-i$ $1$ $-1$
$\theta^{2}$ $1$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$ $1$ $1$
$\theta^{3}$ $1$ $-1$ $1$ $-1$ $-i$ $i$ $-i$ $i$ $1$ $-1$
$X$ $\sqrt{3}$ $\sqrt{3}i$ $-\sqrt{3}$ $-\sqrt{3}i$ $\beta$ $\beta^{3}$ $-\beta$ $-\beta^{3}$ 00 00
$\theta X$ $\sqrt{3}$ $-\sqrt{3}i$ $-\sqrt{3}$ $\sqrt{3}i$ $\beta^{3}$ $\beta$ $-\beta^{3}$ $-\beta$ 00 00
$\theta^{2}X$ $\sqrt{3}$ $\sqrt{3}i$ $-\sqrt{3}$ $-\sqrt{3}i$ $-\beta$ $-\beta^{3}$ $\beta$ $\beta^{3}$ 00 00
$\theta^{3}X$ $\sqrt{3}$ $-\sqrt{3}i$ $-\sqrt{3}$ $\sqrt{3}i$ $-\beta^{3}$ $-\beta$ $\beta^{3}$ $\beta$ 00 00
$Y$ $2$ $2$ $2$ $2$ 00 00 00 00 $-1$ $-1$
$\theta Y$ $2$ $-2$ $2$ $-2$ 00 00 00 00 $-1$ $1$
@@
As in the previous cases, the primaries appearing in the decomposition at the level of characters are the primaries of the extended chiral algebra (i.e. not all Virasoro primaries). Once again there is a Hecke operator relating the parafermion characters to those of the commutant, in this case $\mathsf{T}_{23}$. All lines besides powers of $\theta$ are non-invertible and so give novel defect McKay-Thompson series; $\theta$ implements the global symmetry corresponding to the $\mathbb{Z}_{4A}$ conjugacy class of the monster.
@@ 2.4 $\quantity(D_{3\text{A}},Fi_{23})$: $\mathcal{W}_{D_{3\text{A}}}$ algebra
Finally, we consider the monstralizer pair $\quantity(D_{3\text{A}},Fi_{23})$. In the uplift to VOAs, $D_{3A}$ becomes $\mathcal{W}_{D_{3\text{A}}}$, the chiral algebra of which has been studied in [13, 14, 15]. Comparatively little is known about this VOA: in particular, we are not aware of any attempt to derive the full fusion category of TDLs in the theory. Fortunately, the characters of the theory may be decomposed into minimal model characters and consequently the modular $S$ and $T$ matrices may be derived [6]. We may therefore obtain the Verlinde lines of the theory using Verlinde’s formula [16], but there may be additional TDLs not found by this method. The TDLs found in this way and their eigenvalues are given in table 4.
$h$ 00 $\frac{1}{7}$ $\frac{5}{7}$ $\frac{2}{5}$ $\frac{19}{35}$ $\frac{4}{35}$
$A$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$
@@
$W$ $\varphi$ $\varphi$ $\varphi$ $-\varphi^{-1}$ $-\varphi^{-1}$ $-\varphi^{-1}$
$AW$ $\alpha\varphi$ $\beta\varphi$ $\gamma\varphi$ $-\alpha\varphi^{-1}$ $-\beta\varphi^{-1}$ $-\gamma\varphi^{-1}$
$BW$ $\gamma^{-1}\varphi$ $\alpha^{-1}\varphi$ $\beta^{-1}\varphi$ $-\gamma^{-1}\varphi^{-1}$ $-\alpha^{-1}\varphi^{-1}$ $-\beta^{-1}\varphi^{-1}$
@@
All the Verlinde lines are non-invertible. This is the first case we have considered where the characters of the VOAs forming the monstralizer pair are not known to be related by a Hecke operator; however, the characters of the commutant theory solve a sixth order modular linear differential equation [6], and one may pair them with the correct $\mathcal{W}_{D_{3\text{A}}}$ characters by requiring that
@@ 3 Invariance groups
Now that we have collected a number of defect McKay-Thompson series, we wish to find their modular invariance groups. We do so in two parts: first describing a general method for determining the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and then considering whether the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{R})$ might be larger.
@@ 3.1 In $\operatorname{\textsl{SL}}(2,\mathbb{Z})$
Turning to the invariance groups in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, we begin by exposititing the method used by Lin and Shao to compute the invariance group of the defect McKay-Thompson series found in [5]. As we shall see, this method is not suited to the more general case, and so we develop an algorithm for determining the modular invariance group of any defect McKay-Thompson series in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$.
@@ 3.1.1 Modular $S$ and $T$ matrices
One might hope that Lin and Shao’s procedure could be adapted to find the invariance groups in $SL(2,\mathbb{Z})$ of other defect McKay-Thompson series; to that end, we note that 8 is the lowest value of $n$ for which the modular tranformation $ST^{n}S$ fixes $Z^{\mathcal{D}}(\tau)$ up to a sign. This suggests a general procedure for finding the invariance groups: starting with a defect McKay-Thompson series $Z^{\mathcal{L}}(\tau)$, find the lowest value of $n$ such that
@@ 3.1.2 Coset permutations
where $\rho$ is the representation of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ generated by the modular $S$ and $T$ matrices of $\mathcal{A}$. Moreover, this condition is necessary provided that the characters $\chi_{i}^{(\mathcal{A})}(\tau)$ appearing in the monstralizer decomposition (2.2) are linearly independant. In all the cases under consideration in this work, the various characters either (i) have different leading powers of $q$ or (ii) are equal. In case (i), the characters must obviously be linearly independant. In case (ii) the Hecke images of the redundant characters will also be equal, which means we may represent the McKay-Thompson in an alternate way. Suppose
@@
$\vec{\tilde{\chi}}^{(\mathcal{A})}(\tau)=\matrixquantity(\sqrt{2}\chi_{1}^{(\mathcal{A})}(\tau)\\ \chi_{3}^{(\mathcal{A})}(\tau)\\ \chi_{4}^{(\mathcal{A})}(\tau)\\ ...)\mbox{\quad and\quad}\vec{\tilde{\chi}}^{(\mathcal{B})}(\tau)=\matrixquantity(\sqrt{2}\chi_{1}^{(\mathcal{B})}(\tau)\\ \chi_{3}^{(\mathcal{B})}(\tau)\\ \chi_{4}^{(\mathcal{B})}(\tau)\\ ...).$ (3.6)
@@
where $\tilde{D}_{\mathcal{L}}=\operatorname{\text{diag}}\quantity(\frac{1}{2}(L_{1}+L_{2}),L_{3},L_{4},...)$. We pause to emphasize that equation 3.7 should be understood as a statement about modular functions, not about CFTs. The right hand side of (3.7) should not be interpreted as a “true” defect McKay-Thompson series: we have not constructed a new CFT with characters $\vec{\tilde{\chi}}^{(\mathcal{A})}$. Rather, (3.7) is merely a convenient way of rewriting $\mathcal{Z}^{\mathcal{L}}(\tau)$ in terms of linearly independent objects. Nonetheless, one may check that $\vec{\tilde{\chi}}^{(\mathcal{A})}(\tau)$ and $\vec{\tilde{\chi}}^{(\mathcal{B})}(\tau)$ in fact transform linearly under modular transformations, with “modified” $S$ and $T$-matrices
$\begin{gathered}\tilde{T}^{(\mathcal{A})}=\operatorname{\text{diag}}(T^{(\mathcal{A})}_{11},T^{(\mathcal{A})}_{33},T^{(\mathcal{A})}_{44},...)\\ \tilde{S}^{(\mathcal{A})}_{ij}=\begin{cases}\frac{1}{4}\quantity(S^{(\mathcal{A})}_{11}+S^{(\mathcal{A})}_{12}+S^{(\mathcal{A})}_{21}+S^{(\mathcal{A})}_{22})&i=j=1\\ \frac{1}{2}\quantity(S^{(\mathcal{A})}_{1j}+S^{(\mathcal{A})}_{2j})&i=1,\ j>1\\ \frac{1}{2}\quantity(S^{(\mathcal{A})}_{i1}+S^{(\mathcal{A})}_{i2})&i>1,\ j=1\\ S^{(\mathcal{A})}_{(i+1)(j+1)}&i,j>1\end{cases}\\ \tilde{T}^{(\mathcal{B})}=T^{(\mathcal{A})*}\mbox{\quad and\quad}\tilde{S}^{(\mathcal{B})}=S^{(\mathcal{A})*}.\end{gathered}$ (3.8)
Thus, the problem of finding the modular invariance group of $D_{\mathcal{L}}$ transforming in the representation generated by $S^{(\mathcal{A})}$ and $T^{(\mathcal{A})}$ is identical to that of finding the modular invariance group of $\tilde{D}_{\mathcal{L}}$ transforming in the representation generated by $\tilde{S}^{(\mathcal{A})}$ and $\tilde{T}^{(\mathcal{A})}$. This prodecure allows us to remove the redundant characters, leaving only a linearly independent list behind. Going forward we will have done so, and it may be assumed that the different elements of $\vec{\chi}^{(\mathcal{A})}(\tau)$ and $\vec{\chi}^{(\mathcal{B})}(\tau)$ are linearly independant, in particular having different leading powers of $q$ (though again we note that these elements are no longer necessarily characters of any RCFT, and should be thought of simply as modular functions which we have chosen for convenience).
@@
The utility of this correspondence is in the following fact: a finite-index subgroup of a finitely generated group is entirely determined by the action of the generators on its right cosets, meaning the permutation of the right cosets induced by the action of each generator from the right. This means that, if we know the possible $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$, we can find the invariance group by conjugating each one by $\rho(S)$ and $\rho(T)$, and recording how each generator permutes the conjugates. To see this, begin with $D_{\mathcal{L}}$ and successively conjugate with $\rho(S)$ and $\rho(T)$, building a directed graph with two edges exiting each node, until the graph closes on itself. This is the Cayley graph corresponding to the action of $SL(2,\mathbb{Z})$ on the cosets; the graph will contain as many nodes as there are distinct $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$ (which is the same as the number of cosets of $\Gamma$, a.k.a. the index of $\Gamma$ in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$). Then, find a generating set for the fundamental group of the directed graph. Each of the elements of the generating set is a sequence of edges labelled with generators of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ and so may be translated into an element of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$; these elements then generate $\Gamma$. In practice, once we have built the Cayley graph, we may number the nodes arbitrarily and find the permutation of cosets induced by $S$ (resp. $T$) by following the chain of consecutive $S$ (resp. $T$) edges. Given the permutations, the Sage mathematics system provides a built-in method for determining the invariance group [18]. An example graph is shown in figure 1.
1 2 3 5 6 7 8 4 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 $S$$T$

### v2

@@ 11 $\sigma$ $\epsilon$
Our next pair of monstralizer VOAs are the chiral algebra of the $\mathbb{Z}_{4}$ parafermion CFT and its commutant in $V^{\natural}$. Similar to the $\mathbb{Z}_{3A}$ case, we may take the monstralizing RCFT to be the $\mathbb{Z}_{4}$ parafermion CFT. As in the $\mathbb{Z}_{3}$ case, the $\mathbb{Z}_{4}$ parafermion CFT contains topological defects which commute with Virasoro but not the fully extended chiral algebra (in fact, infinitely many). However, to find the defect McKay-Thompson series associated to these, our method would require us to decompose the modular invariant of the $\mathbb{Z}_{4}$ parafermion CFT into finitely many Virasoro characters. As the theory is irrational under Virasoro, this is not possible, so we satisfy ourselves with TDLs preserving the fully extended chiral algebra (under which the theory is rational). These lines are given in table 3 [13].
$(l,m)$ $(4,4)$ $(4,2)$ $(4,0)$ (4,2)(4,-2) $(3,3)$ $(3,1)$ (3,1)(3,-1) $(1,1)$ $(2,0)$ $(2,2)$
11 11 11 11 11 11 11 11 11 11 11
$\theta$ 11 1-1 11 1-1 $i$ i-i $i$ i-i 11 1-1
$\theta^{2}$ 11 11 11 11 1-1 1-1 1-1 1-1 11 11
$\theta^{3}$ 11 1-1 11 1-1 i-i $i$ i-i $i$ 11 1-1
$X$ $\sqrt{3}$ $\sqrt{3}i$ $\sqrt{3}$ $\sqrt{3}i$ $\beta$ $\beta^{3}$ $\beta$ $\beta^{3}$ 0 0
$\theta X$ $\sqrt{3}$ $\sqrt{3}i$ $\sqrt{3}$ $\sqrt{3}i$ $\beta^{3}$ $\beta$ $\beta^{3}$ $\beta$ 0 0
$\theta^{2}X$ $\sqrt{3}$ $\sqrt{3}i$ $\sqrt{3}$ $\sqrt{3}i$ $\beta$ $\beta^{3}$ $\beta$ $\beta^{3}$ 0 0
$\theta^{3}X$ $\sqrt{3}$ $\sqrt{3}i$ $\sqrt{3}$ $\sqrt{3}i$ $\beta^{3}$ $\beta$ $\beta^{3}$ $\beta$ 0 0
$Y$ 22 22 22 22 0 0 0 0 1-1 1-1
$\theta Y$ 22 2-2 22 2-2 0 0 0 0 1-1 11
@@
Once again there is a Hecke operator relating the parafermion characters to those of the commutant, in this case $\mathsf{T}_{23}$. All lines besides powers of $\theta$ are non-invertible and so give novel defect McKay-Thompson series; $\theta$ implements the global symmetry corresponding to the $\mathbb{Z}_{4A}$ conjugacy class of the monster.
@@ 2.4 $\quantity(D_{3\text{A}},Fi_{23})$: $\mathcal{W}_{D_{3\text{A}}}$ algebra
Finally, we consider the monstralizer pair $\quantity(D_{3\text{A}},Fi_{23})$. In the uplift to VOAs, $D_{3A}$ becomes the chiral algebra $\mathcal{W}_{D_{3\text{A}}}$, which has been studied in [15, 16, 17]. We take the monstralizing RCFT to be the theory with full chiral algebra $\mathcal{W}_{D_{3\text{A}}}\otimes\overline{\mathcal{W}}_{D_{3\text{A}}}$ and the diagonal modular invariant. Comparatively little is known about this VOA: in particular, we are not aware of any attempt to derive the full fusion category of TDLs in the theory. Fortunately, the characters of the theory may be decomposed into minimal model characters and, consequently, the modular $S$ and $T$ matrices may be derived [6]. We may therefore obtain the Verlinde lines of the theory using Verlinde’s formula [18], but there may be additional TDLs (preserving Virasoro but not the fully extended chiral algebra) not found by this method. The TDLs found in this way and their eigenvalues are given in table 4.
$h$ 0 $\frac{1}{7}$ $\frac{5}{7}$ $\frac{2}{5}$ $\frac{19}{35}$ $\frac{4}{35}$
11 11 11 11 11 11 11
AA $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$
@@
$W$ $\varphi$ $\varphi$ $\varphi$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$
AWAW $\alpha\varphi$ $\beta\varphi$ $\gamma\varphi$ $\alpha\varphi^{-1}$ $\beta\varphi^{-1}$ $\gamma\varphi^{-1}$
BWBW $\gamma^{-1}\varphi$ $\alpha^{-1}\varphi$ $\beta^{-1}\varphi$ $\gamma^{-1}\varphi^{-1}$ $\alpha^{-1}\varphi^{-1}$ $\beta^{-1}\varphi^{-1}$
@@
All the Verlinde lines are non-invertible. This is the first case we have considered where the characters of the VOAs forming the monstralizer pair are not known to be related by a Hecke operator; however, the characters of the commutant theory solve a sixth-order modular linear differential equation [6], and one may pair them with the correct $\mathcal{W}_{D_{3\text{A}}}$ characters by requiring that
@@ 3 Invariance groups
Now that we have collected a number of defect McKay-Thompson series, we wish to find their modular invariance groups. We do so in two parts: first, we describe a general method to determine the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and then consider whether the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{R})$ might be larger.
@@ 3.1 In $\operatorname{\textsl{SL}}(2,\mathbb{Z})$
Turning to the invariance groups in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, we begin by expositing the method used by Lin and Shao to compute the invariance group of the defect McKay-Thompson series found in [5]. As we shall see, this method is not suited to the more general case, and so we develop an algorithm to determine the modular invariance group of any defect McKay-Thompson series in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$.
@@ 3.1.1 Modular $S$ and $T$ matrices
One might hope that Lin and Shao’s procedure could be adapted to find the invariance groups in $SL(2,\mathbb{Z})$ of other defect McKay-Thompson series; to that end, we note that 8 is the lowest value of $n$ for which the modular transformation $ST^{n}S$ fixes $Z^{\mathcal{D}}(\tau)$ up to a sign. This suggests a general procedure for finding the invariance groups: starting with a defect McKay-Thompson series $Z^{\mathcal{L}}(\tau)$, find the lowest value of $n$ such that
@@ 3.1.2 Coset permutations
where $\rho$ is the representation of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ generated by the modular $S$ and $T$ matrices of $\mathcal{A}$. Moreover, this condition is necessary provided that the characters $\chi_{i}^{(\mathcal{A})}(\tau)$ appearing in the monstralizer decomposition (2.1) are linearly independent. In all the cases under consideration in this work, the various characters either (i) have different leading powers of $q$ or (ii) are equal. In case (i), the characters must obviously be linearly independent. In case (ii) the Hecke images of the redundant characters will also be equal, which means we may represent the McKay-Thompson in an alternate way. Suppose that
@@
[…]
@@
where
$\quantity(\tilde{D}_{\mathcal{L}})_{ij}=\begin{cases}\frac{1}{2}\quantity(\quantity(D_{\mathcal{L}})_{11}+\quantity(D_{\mathcal{L}})_{12}+\quantity(D_{\mathcal{L}})_{21}+\quantity(D_{\mathcal{L}})_{22})&i=j=1\\ \frac{1}{\sqrt{2}}\quantity(\quantity(D_{\mathcal{L}})_{1(j+1)}+\quantity(D_{\mathcal{L}})_{2(j+1)})&i=1,\ j>1\\ \frac{1}{\sqrt{2}}\quantity(\quantity(D_{\mathcal{L}})_{(i+1)1}+\quantity(D_{\mathcal{L}})_{(i+1)2})&i>1,\ j=1\\ \quantity(D_{\mathcal{L}})_{(i+1)(j+1)}&i,j>1\end{cases}.$ (3.8)
We pause to emphasize that equation 3.7 should be understood as a statement about modular functions, not about CFTs. The right-hand side of (3.7) does not correspond to a true decomposition into subVOA representations: we have not constructed a new CFT with characters $\vec{\tilde{\chi}}^{(\mathcal{A})}$. Rather, (3.7) is merely a convenient way of rewriting $\mathcal{Z}^{\mathcal{L}}(\tau)$ in terms of linearly independent objects. Nonetheless, one may check that $\vec{\tilde{\chi}}^{(\mathcal{A})}(\tau)$ and $\vec{\tilde{\chi}}^{(\mathcal{B})}(\tau)$ in fact transform linearly under modular transformations, with “modified” $S$ and $T$-matrices
$\begin{gathered}\tilde{T}^{(\mathcal{A})}=\operatorname{\text{diag}}(T^{(\mathcal{A})}_{11},T^{(\mathcal{A})}_{33},T^{(\mathcal{A})}_{44},...)\\ \tilde{S}^{(\mathcal{A})}_{ij}=\begin{cases}\frac{1}{4}\quantity(S^{(\mathcal{A})}_{11}+S^{(\mathcal{A})}_{12}+S^{(\mathcal{A})}_{21}+S^{(\mathcal{A})}_{22})&i=j=1\\ \frac{1}{2}\quantity(S^{(\mathcal{A})}_{1(j+1)}+S^{(\mathcal{A})}_{2(j+1)})&i=1,\ j>1\\ \frac{1}{2}\quantity(S^{(\mathcal{A})}_{(i+1)1}+S^{(\mathcal{A})}_{(i+1)2})&i>1,\ j=1\\ S^{(\mathcal{A})}_{(i+1)(j+1)}&i,j>1\end{cases}\\ \tilde{T}^{(\mathcal{B})}=T^{(\mathcal{A})*}\mbox{\quad and\quad}\tilde{S}^{(\mathcal{B})}=S^{(\mathcal{A})*}.\end{gathered}$ (3.9)
Thus, the problem of finding the modular invariance group of $D_{\mathcal{L}}$ transforming in the representation generated by $S^{(\mathcal{A})}$ and $T^{(\mathcal{A})}$ is identical to that of finding the modular invariance group of $\tilde{D}_{\mathcal{L}}$ transforming in the representation generated by $\tilde{S}^{(\mathcal{A})}$ and $\tilde{T}^{(\mathcal{A})}$. This procedure allows us to remove the redundant characters, leaving only a linearly independent list behind. Going forward we will have done so, and it may be assumed that the different elements of $\vec{\chi}^{(\mathcal{A})}(\tau)$ and $\vec{\chi}^{(\mathcal{B})}(\tau)$ are linearly independent, in particular having different leading powers of $q$ (though again we note that these elements are no longer necessarily characters of any RCFT, and should be thought of simply as modular functions which we have chosen for convenience).
@@
The utility of this correspondence is in the following fact: a finite-index subgroup of a finitely generated group is entirely determined by the action of the generators on its right cosets, meaning the permutation of the right cosets induced by the action of each generator from the right. This means that, if we know the possible $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$, we can find the invariance group by conjugating each one by $\rho(S)$ and $\rho(T)$, and recording how each generator permutes the conjugates. To see this, begin with $D_{\mathcal{L}}$ and successively conjugate with $\rho(S)$ and $\rho(T)$, building a directed graph with two edges exiting each node, until the graph closes on itself. This is the Cayley graph corresponding to the action of $SL(2,\mathbb{Z})$ on the cosets; the graph will contain as many nodes as there are distinct $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$ (which is the same as the number of cosets of $\Gamma$, a.k.a. the index of $\Gamma$ in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$). Then, find a generating set for the fundamental group of the directed graph. Each of the elements of the generating set is a sequence of edges labeled with generators of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ and so may be translated into an element of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$; these elements then generate $\Gamma$. In practice, once we have built the Cayley graph, we may number the nodes arbitrarily and find the permutation of cosets induced by $S$ (resp. $T$) by following the chain of consecutive $S$ (resp. $T$) edges. Given the permutations, the Sage mathematics system provides a built-in method for determining the invariance group [20]. An example graph is shown in figure 1.
1 2 3 5 6 7 8 4 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 SSTT